OpenStudy (anonymous):

If Dr. Suzuki rides his bike to his office, he averages 6 miles per hour. If he drives his car, he averages 12 miles per hour. His time driving is 1/2 hour less than his time bicycling. How far is his office?

OpenStudy (mathmate):

D/6mph - D/12mph = 1/2 hr (2D-D) = (1/2)*12 D=6 miles